OpenStudy (anonymous):

Given this function, find the domain and range, the x and y intercepts, the horizontal asymptote, and the vertical asymptote.

OpenStudy (anonymous):

\[f(x)= \frac{ 2 }{ ^{x^2 - 2x - 3} }\]

OpenStudy (anonymous):

range is \[x \in R - \left\{ 2 \right\}\left\{ 3 \right\}\]

OpenStudy (anonymous):

domain is infinity

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you help me with the other things?

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

which ones

OpenStudy (anonymous):

The x and y intercepts, the horizontal asymptote, and the vertical asymptote

OpenStudy (anonymous):

horizontal asymptote put x=0 vertical asymptote put y=0

OpenStudy (anonymous):

actually those answers are not correct

OpenStudy (anonymous):

in fact none of them are correct

OpenStudy (anonymous):

you know midnight study leads to this :D sry

OpenStudy (anonymous):

for the DOMAIN set the denominator equal to zero and solve then exclude those values \[x^2-2x-3=0\] \[(x-3)(x+1)=0\] \[x=3,x=-1\] so the domain is all numbers except \(3\) and \(-1\)

OpenStudy (anonymous):

range is all real numbers except 0 because a fraction is only 0 if the numerators is 0, and your numerator is 2

OpenStudy (anonymous):

vertical asymptote you already solved for, it is the zeros of the denominator \[x=3,x=-1\] horizontal asymptote is \(y=0\) because the degree of the numerator is less than the degree of the denominator

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Yes, I got that too :)

OpenStudy (aravindg):

x intercept is when y=0